# Bulk–boundary correspondence from hyper-invariant tensor networks

### v1

@@ Appendix C On bipartite unitary matrices
It turns out jonnadula2017impact that the entire family of qubit dual unitary matrices (17) can be characterized by a single parameter $\alpha_{3}\in[0,\pi/4]$, with $\alpha_{1}=\alpha_{2}=\pi/4$. Thus they lie on the edge interpolating between $DCNOT$ and SWAP gates. An alternative parametrization of this family, up to local unitaries, is presented in (13).
The Weyl chamber is more than just a nice way to visualize the set of bipartite unitary gates, it captures its geometry as well. For instance, the work Dual_u_max_dist proves that the dual unitary gates in any dimension maximize the distance from the product of local gates, represented as $Id$ vertex.
Another worth-mentioning property of dual unitary matrices is their operator Schmidt decomposition into products of matrices acting on both subsystems separately PhysRevA.67.052301 . Unitarity condition for the matrix with reshuffled indices (17, right) implies Dual_u_max_dist that all the Schmidt coefficients are equal. This fact assures the maximal entanglement entropy in the space of all unitary matrices in $U(d^{2})$, which can be interpreted that the dual unitary form the set of ’maximally non-local’ bipartite gates.
Appendix D Decay of two-point correlation function
@@
$M^{n}\approx\frac{1}{D}(\sum_{\mathbb{i}}|\mathbb{i}\mathbb{i}\rangle)(\sum_{\mathbb{i}}\langle\mathbb{i}\mathbb{i}|)+\tbinom{n}{k-1}\lambda_{2}^{n-(k-1)}|\mathbb{v_{R}}\rangle\langle\mathbb{v_{L}}|,$ (22)
where $\langle\mathbb{v_{L}}|$ and $|\mathbb{v_{R}}\rangle$ are appropriate row and column of $\beta^{-1}$ and $\beta$ respectively, and can be interpreted as generalized eigenvectors of $M$ to the eigenvalue $\lambda_{2}$. Therefore, by the property (21), we effectively recreated the situation discussed in Corollary 1.
@@
Appendix E Numerical calculation of three-point correlation functions
@@
$N:=\frac{\left(M-\frac{1}{D}(\sum_{\mathbb{i}}|\mathbb{i}\mathbb{i}\rangle)(\sum_{\mathbb{i}}\langle\mathbb{i}\mathbb{i}|)\right)^{n}}{\lambda_{2}^{n}}\approx\tbinom{n}{k-1}|\mathbb{v_{R}}\rangle\langle\mathbb{v_{L}}|+O\left(n^{-1}\right)~{},$ (23)
@@
$N=U\Sigma V$, with the row of $V$ and column of $U$ corresponding to the largest singular value approximating $\langle\mathbb{v_{L}}|$ and $|\mathbb{v_{R}}\rangle$. In practice, we chose $n=20$ and dismissed the cases in which the ratio between the leading and next to leading singular value of $N$ were smaller than $100$ to ensure the accuracy of the approximation.
@@
Figure 10: Violin plots presenting values of contractions of inner tensors with the bulk operator as in Figure 6(b). (a) The distribution obtained from $10^{5}$ random bulk operators and random tensor building blocks, drawn as in Figure 8. (b) The distribution obtained from $10^{5}$ samples with tensor constructed form unitaries (13), (14) with $a=0.302$, $b=0.817$ corresponding to minimal $\Delta$ in Figure 7(c). In each case, we draw the bulk operators from Hermitian, positive semi-definite matrices with trace $1$ and considered triples of appropriate left normalized generalized eigenvectors $\langle\mathbb{v_{L}}|$, as discussed above. The maximal, minimal and median values are highlighted.
We note that the approximations of both $|\mathbb{v_{R}}\rangle$ and $\langle\mathbb{v_{L}}|$ by SVD are normalized, which is not the case in Jordan decomposition. Thus we lost the proportionality factor which cannot be re-obtained without a knowledge of the size $k$ of the largest Jordan block corresponding to $\lambda_{2}$. However, since there can be more than one block to this eigenvalue, one cannot determine $k$ by simply calculating the multiplicity of generalized eigenvectors of $M$. Thus we return to the problem of performing Jordan decomposition numerically.
@@
be Hermitian, positive semi-definite with trace $1$. In each scenario, we considered triples of appropriate left normalized generalized eigenvectors $\langle\mathbb{v_{L}}|$. We note that resulting values are real up to numerical precision, which is in accordance with the fact that the original quantity, expectation value for a mapped hermitian operator, must be real as well.

### v2

@@ V Concluding Remarks
It turns out [41] that the entire family of qubit dual unitary matrices (17) can be characterized by a single parameter $\alpha_{3}\in[0,\pi/4]$, with $\alpha_{1}=\alpha_{2}=\pi/4$. Thus they lie on the edge interpolating between $DCNOT$ and SWAP gates. An alternative parametrization of this family, up to local unitaries, is presented in (13).
The Weyl chamber is more than just a nice way to visualize the set of bipartite unitary gates, it captures its geometry as well. For instance, the work [42] proves that the dual unitary gates in any dimension maximize the distance from the product of local gates, represented as $Id$ vertex.
Another worth-mentioning property of dual unitary matrices is their operator Schmidt decomposition into products of matrices acting on both subsystems separately [43]. Unitarity condition for the matrix with reshuffled indices (17, right) implies [42] that all the Schmidt coefficients are equal. This fact assures the maximal entanglement entropy in the space of all unitary matrices in $U(d^{2})$, which can be interpreted that the dual unitary form the set of ’maximally non-local’ bipartite gates.
Appendix D Appendix D. Decay of two-point correlation function
@@
$M^{n}\approx\frac{1}{D}(\sum_{\mathbf{i}}|\mathbf{i}\mathbf{i}\rangle)(\sum_{\mathbf{i}}\langle\mathbf{i}\mathbf{i}|)+\tbinom{n}{k-1}\lambda_{2}^{n-(k-1)}|\mathbf{v_{R}}\rangle\langle\mathbf{v_{L}}|,$ (23)
where $\langle\mathbf{v_{L}}|$ and $|\mathbf{v_{R}}\rangle$ are appropriate row and column of $\beta^{-1}$ and $\beta$ respectively, and can be interpreted as generalized eigenvectors of $M$ to the eigenvalue $\lambda_{2}$. Therefore, by the property (22), we effectively recreated the situation discussed in Corollary 1.
@@
Appendix E Appendix E. Numerical calculation of three-point correlation functions
@@
$N:=\frac{\left(M-\frac{1}{D}(\sum_{\mathbf{i}}|\mathbf{i}\mathbf{i}\rangle)(\sum_{\mathbf{i}}\langle\mathbf{i}\mathbf{i}|)\right)^{n}}{\lambda_{2}^{n}}\approx\tbinom{n}{k-1}|\mathbf{v_{R}}\rangle\langle\mathbf{v_{L}}|+O\left(n^{-1}\right)~{},$ (24)
@@
$N=U\Sigma V$, with the row of $V$ and column of $U$ corresponding to the largest singular value approximating $\langle\mathbf{v_{L}}|$ and $|\mathbf{v_{R}}\rangle$. In practice, we chose $n=20$ and dismissed the cases in which the ratio between the leading and next to leading singular value of $N$ were smaller than $100$ to ensure the accuracy of the approximation.
@@
Figure 12: Violin plots presenting values of contractions of inner tensors with the bulk operator as in Figure 7(b). (a) The distribution obtained from $10^{5}$ random bulk operators and random tensor building blocks, drawn as in Figure 10. (b) The distribution obtained from $10^{5}$ samples with tensor constructed form unitaries (13), (14) with $a=0.302$, $b=0.817$ corresponding to minimal $\Delta$ in Figure 9(c). In each case, we draw the bulk operators from Hermitian, positive semi-definite matrices with trace $1$ and considered triples of appropriate left normalized generalized eigenvectors $\langle\mathbf{v_{L}}|$, as discussed above. The maximal, minimal and median values are highlighted.
We note that the approximations of both $|\mathbf{v_{R}}\rangle$ and $\langle\mathbf{v_{L}}|$ by SVD are normalized, which is not the case in Jordan decomposition. Thus we lost the proportionality factor which cannot be re-obtained without a knowledge of the size $k$ of the largest Jordan block corresponding to $\lambda_{2}$. However, since there can be more than one block to this eigenvalue, one cannot determine $k$ by simply calculating the multiplicity of generalized eigenvectors of $M$. Thus we return to the problem of performing Jordan decomposition numerically.
@@
be Hermitian, positive semi-definite with trace $1$. In each scenario, we considered triples of appropriate left normalized generalized eigenvectors $\langle\mathbf{v_{L}}|$. We note that resulting values are real up to numerical precision, which is in accordance with the fact that the original quantity, expectation value for a mapped hermitian operator, must be real as well.